OpenStudy (anonymous):

“By Completing the square turn your expression, given in expanded form, into Turning point form and Sketch the graph using equation in Turning point form”. x^2 + 6x +2

OpenStudy (anonymous):

dont need to sketch the graph i can do so

OpenStudy (solomonzelman):

not sure what you need, but to make the `x²+6x+2` into a perfect square, you would need to add 7 to it.

OpenStudy (solomonzelman):

not sure what you need, but to make the `x²+6x+2` into a perfect square, you would need to add 7 to it.

OpenStudy (anonymous):

why 7? :)

OpenStudy (solomonzelman):

because (looking at `ax²+bx+c`) when `a=1`, you need the `c` to be equal to `(b/2)²`.

OpenStudy (solomonzelman):

you will get `(x+3)²` for that part.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{y= x^2 + 6x +2 \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{y= x^2 + 6x +2 \color{darkgoldenrod}{+7} \color{red}{-7} \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{y= x^2 + 6x +2 +7 \color{red}{-7} \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{y= x^2 + 6x +9\color{red}{-7} \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{y= (x+3)^2\color{red}{-7} \LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{y= (x+3)^2-7 \LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

center (-3,-7)

OpenStudy (anonymous):

I still don't quite understand where the 7 is coming from?

OpenStudy (anonymous):

and also is the center the turning point?

OpenStudy (solomonzelman):

For the number of 7, you see what I did... I added 7 and subtracted 7 without canceling them out. Then you see that I used to positive 7 for the quadratic to get the perfect square quadratic, leaving the -7 alone.... And yes, the turning point is the center.

OpenStudy (solomonzelman):

Which we found to be (-3,-7)

OpenStudy (anonymous):

but where does the 7 come from? are you just substituting it in as a random number? :)